OpenStudy (anonymous):

If .6 grams of magnesium is in 1 gram of magnesium oxide, how much magnesium is in .84 grams?

OpenStudy (anonymous):

hlppp

OpenStudy (anonymous):

this is writing.... you might want to post this again in Mathematics

OpenStudy (texaschic101):

proportion time.. 0.6 / 1 = x / 0.84 --- 0.6 grams of m to 1 gram of mo = x grams of m to 0.84 grams of mo. cross multiply (1)(x) = (0.6)(0.84) x = 0.504 grams so...if there is 0.84 grams of magnesium oxide, then there is 0.504 grams of magnesium used.

OpenStudy (nevermind_justschool):

Please don't forget to close the question now that you got help. :)